# Semi-orthogonality in Fukaya-Seidel mirrors to blowups of abelian varieties

### v1

@@ 1.2. Statement of main theorem
There exist two full subcategories $\mathcal{F}_{1}$ (U-shaped Lagrangians) and $\mathcal{F}_{2}$ (Lagrangians fibered over cotangent fibers) of the Fukaya category $H^{*}FS(Y^{0},W_{0})$ which are equivalent to $D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$ and $D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$, respectively. Furthermore, for the full subcategory $\mathcal{F}:=\mathcal{F}_{1}\cup\mathcal{F}_{2}$, there is an equivalence
(1.7) $\mathcal{F}\xrightarrow[]{\cong}D^{b}_{\mathcal{L}}\mathrm{Coh}(X).$
@@
The wrapping phenomenon that appears when we puncture $Y$ to $Y^{0}$ also appears in related settings such as [GPS24, §3.3] where stop removal corresponds to localization.
@@
For a suitable definition of $FS(Y^{0},W_{0})$, upgrading the theorem to an $A_{\infty}$-result, and for a suitable adaptation of core HMS [PS23], we expect this theorem can be upgraded to the originally conjectured equivalence in Equation (1.1) which would be $D^{b}Coh(X)\cong D^{\pi}FS(Y^{0},W_{0})$.
@@
We directly construct Lagrangians and calculate their morphisms in $(Y^{0},W_{0})$. Another approach could be to combine the left diagram of Figure 1 for $H^{0}FS(Y,v_{0})$ from [Can20], prove a similar diagram for $H^{0}FS(V^{\vee}\times\mathbb{C}^{*}_{y^{\vee}},y^{\vee})$ as in the right diagram of Figure 1, and then combine them as suggested in [Aur18, Remark 1.5] and also discussed in [Syl19, §3.3: Gluing].
$D^{b}_{\mathcal{L}}\mathrm{Coh}(V)$$D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$$H^{0}\mathrm{Fuk}(V^{\vee})$$H^{0}\mathrm{FS}(Y,v_{0})$$\scriptstyle{\iota^{*}}$$\scriptstyle{\text{HMS on }V}$$\scriptstyle{\text{HMS on }H=\Sigma_{2}}\scriptstyle{\cup}$
$D^{b}_{\mathcal{L}}\mathrm{Coh}(V)$$D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$$H^{0}\mathrm{Fuk}(V^{\vee})$$H^{0}\mathrm{FS}(V^{\vee}\times\mathbb{C}^{*}_{y^{\vee}},y^{\vee})$$\scriptstyle{\mathrm{pr}_{V}^{*}}$$\scriptstyle{\text{HMS on }V}$
@@
Outline of the paper. In Section 2 we describe the B-side semi-orthogonality of $D^{b}\text{Coh}(X)$. In Section 3, we define the mirror Lagrangian objects in Subsection 3.1 and morphisms in the Fukaya category of $(Y^{0},W_{0})$ in Subsection 3.2. This uses standard theory where the details to fill in for this setting are to define the Lagrangians, continuation maps, and prove the latter are well-defined. We define how Lagrangians wrap in Subsection 3.2.1, and give an overview of how to pass from the cochain complex to the localized category in 3.2.2. Then in Subsections 3.2.3 and 3.2.4 we define the quasi-units we localize at, and their inverses, the continuation maps. Finally in Subsection 3.2.5 we define the morphisms by taking a homotopy colimit over multiplication by quasi-units, using a model from [AA24] that amounts to wrapping one way on the input Lagrangian and in the opposite direction on the output in Equation (3.29). In Subsection 3.2.6 we carry out the calculations of the morphisms in Lemmas 3.20 (within $\mathcal{F}_{1}$), 3.23 (nonzero from $\mathcal{F}_{1}$ to $\mathcal{F}_{2}$), 3.24 (zero from $\mathcal{F}_{2}$ to $\mathcal{F}_{1}$), and 3.25 (within $\mathcal{F}_{2}$). We then move on to the product structure, to prove the HMS map is indeed a functor, in Section 4. This requires an analysis of gradings to determine what intersection points appear in the product, which is Lemma 4.3. We then prove the main theorem, Theorem 1.8, in Section 5.
Acknowledgements. We thank Denis Auroux for helpful comments in formulating the direction of this project and for explaining his joint work [AA24]. We thank Hiro Lee Tanaka, Sheel Ganatra, and Zack Sylvan for helpful discussions on wrapping in Fukaya categories, Christian Schnell and Alexander Kuznetsov for helpful discussions on semi-orthogonal decompositions of $D^{b}\text{Coh}$, and Katrin Wehrheim for helpful discussions on moduli spaces. Thanks to Haniya Azam, Hiro Lee Tanaka, and Sheel Ganatra for helpful comments on an earlier draft. C. Cannizzo was partially supported by NSF DMS-2316538. She also thanks the Collaborative Research Centre TRR 326 Geometry and Arithmetic of Uniformized Structures (GAUS) at Goethe Universität-Frankfurt for hosting her during the completion of this work.
@@ 2. B-side semi-orthogonal decomposition
2.1. Orthogonality: morphisms from $H$ to $V\times\mathbb{C}$
@@
(2.1) $\operatorname{Ext}(\pi^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}}),j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{p}(-1)))=0$
which we will see matches with the A-side. We also need to show that morphisms in the other direction match with the A-side, so we consider those next.
2.2. Morphisms in the other direction: $V\times\mathbb{C}$ to $H$
By the adjunction formula $K_{X}=\pi^{*}K_{V\times\mathbb{C}}\otimes\mathcal{O}(E)$. As in [Huy05, Proposition 2.5.5], we define the relative canonical bundle as follows,
@@ 2.2. Morphisms in the other direction: $V\times\mathbb{C}$ to $H$
where $\mathcal{O}_{E}(E):=j^{*}\mathcal{O}(E)=\mathcal{O}_{p}(-1)$. Furthermore, $j^{!}$ admits a left adjoint $j_{*}$ for $j:E\to X$ where
@@
Note that $\otimes\mathcal{O}_{E}(E)$ is an auto-equivalence of $D^{b}\text{Coh}(X)$, thus its contribution cancels. Also, the statement that Equation (1.5) commutes is that $\pi\circ j=i\circ p$. Thirdly, $p:E\to H\times\{0\}$ is surjective so $p^{*}$ is fully faithful, namely, it is an isomorphism on hom sets. Hence we can simplify to obtain
@@
3. A-side semi-orthogonality
We define the objects and morphisms in the wrapped derived Fukaya category for the symplectic Landau-Ginzburg model $(Y^{0},W_{0})$ with symplectic form $\omega$ given by that of [Can20, §3.5] restricted to $Y^{0}$. This builds on the constructions of many Fukaya categories: the original $A_{\infty}$-category [FOOO09] and Fukaya-Seidel category for Lefschetz fibrations [Sei08], as well as [AS10, GPS20] for exact symplectic manifolds, and [AA24] for Landau-Ginzburg models, to name a few.
The Fukaya category roughly consists of Lagrangian submanifolds as objects and cochain complexes generated by their intersection points as morphisms, graded by the Maslov index. The differential counts bigons between input and output intersection points, weighted by area. An account of the Fukaya category for arbitrary symplectic form, including non-exact and with B-field, as well as how structure maps change under Lagrangian isotopy, is given in [ACLL23a]. Here we take the B-field to be zero. For gradings in a general symplectic Landau-Ginzburg model including non-exact and non-Lefschetz, we use [ACLL23b, §3].
@@ 3. A-side semi-orthogonality
Since $Y^{0}$ is non-compact, we take the Fukaya-Seidel category on $W_{0}:Y^{0}\to\mathbb{C}$. To use the machinery for Fukaya-Seidel categories, we need to define the continuation maps and show they are well-defined by analyzing the geometry of $(Y^{0},W_{0})$. On the punctured Landau-Ginzburg model, the projection of Lagrangians in $\mathcal{F}_{2}$ under $W_{0}$ is a curve that can fully wrap in the base around the puncture. All projected Lagrangians still partially wrap at infinity, as is defined for Lagrangians in a Landau-Ginzburg model to $\mathbb{C}$ [AS19, Appendix A], [GPS20, §1.2, Example 1], [AA24]. Since fibers and their Lagrangians are compact, we do not wrap Lagrangians in the fibers. These features give the definition of our continuation maps in Definition 3.9.
Remark 3.1 (Two pictorial representations for the base: $\mathbb{C}^{*}$ and cylinder).
@@
The base of $W_{0}$ is topologically a cylinder $\mathbb{C}^{*}$, and a generic fiber is $T^{4}$. Curves drawn in a plane represent projections of Lagrangians under $W_{0}$, while curves drawn on a cylinder represent projections of Lagrangians under $W_{0}$ composed with $\mathbb{C}^{*}\to\mathbb{R}\times S^{1}$ given by $re^{2\pi i\theta}\mapsto(r,\theta)$.
@@
3.1. Objects in $\mathcal{F}_{1}$, $\mathcal{F}_{2}$
Lagrangians are “thimble-like” in the sense of [GPS24, Definition 8.19], as follows.
Definition 3.2 (Parallel transport, [MS98, Equation (6.2)]).
Let $\Phi_{\gamma(0)\to\gamma(t)}:W_{0}^{-1}(\gamma(0))\to W_{0}^{-1}(\gamma(t))$ denote parallel transport in $(Y^{0},W_{0})$ with respect to $\omega|_{Y^{0}}$ (for $\omega$ on $Y$ defined in [Can20, §3.5]) over an embedded path $\gamma:[0,t]\to\mathbb{C}^{*}$ avoiding the critical value $\text{critv}(W_{0})=-T^{\epsilon}$.
We define two subcategories $\mathcal{F}_{1},\mathcal{F}_{2}$ of the Fukaya-Seidel category of $(Y^{0},W_{0})$. Define a point
@@ 3.1. Objects in $\mathcal{F}_{1}$, $\mathcal{F}_{2}$
where $(\theta_{1},\theta_{2})=\frac{1}{2\pi}(\arg(t_{1}),\arg(t_{2}))$ for $(t_{1},t_{2})\in(\mathbb{C}^{*})^{2}$ the affine coordinates corresponding to monomials $\chi^{(1,0,0)}$ and $\chi^{(0,1,0)}$ on the polytope $\Delta_{Y}$ of Equation (1.3).
Definition 3.4 (Objects of $\mathcal{F}_{1}$).
Objects of $\mathcal{F}_{1}$ are $\{L_{j}[k]\}_{j,k\in\mathbb{Z}}$ defined as
@@
such that
$r(0)=S,\qquad\theta(0)=-\frac{1}{2},\qquad r^{\prime}(t)=1\;\forall t>>0\text{ and }\forall t<<0,\qquad\theta(t)=\begin{cases}\theta_{h_{1}}(L_{j})&\forall t>0\,s.t.~r(t)>R_{4}\\ \theta_{h_{2}}(L_{j})&\forall t<0\,s.t.~r(t)>R_{4}\end{cases},$
for $R_{4}$ as in Equation (3.3) and where $r^{\prime}(t)=1$ by the time $r(t)$ reaches $R_{4}$, so that $\gamma_{L_{j}}$ is a U-shape in the base of $W_{0}$ around $T^{\epsilon}$ whose ends are angular lines at constant angles $\theta_{h_{1}}(L_{j})<\theta_{h_{2}}(L_{j})$ in the interval (1/2,1/2)(-1/2,1/2) on each end. In particular, $\gamma_{L_{j}}(0)=-S$.
@@
U-shaped Lagrangians have appeared in the literature such as [AS19, §A.2], [ACLL23b], and [AA24]. Here we use the model that ends of the U-shape project to angular rays away from 0.
Definition 3.6 (Objects of $\mathcal{F}_{2}$).
Objects of $\mathcal{F}_{2}$ are $\{K_{j}[k]\}_{j,k\in\mathbb{Z}}$ defined as
@@
such that
$r(0)=S,\qquad r(1)=0,\qquad\theta(0)=-\frac{1}{2},\qquad(r(t),\theta(t))=(t,\theta_{h}(K_{j})),\;\forall t>1$
so that $\delta_{K_{j}}$ is an arc from S-S, passing through 0, and is then a single arc at fixed angle $\theta_{h}(K_{j})$. To be the correct mirror, we take the angle to lie above any U-shape ends so $\theta_{h}(K_{j})>\theta_{h^{\prime}}(L_{j^{\prime}})$ in (1/2,1/2)(-1/2,1/2) $\forall j^{\prime}\in\mathbb{Z}$.
Parallel transport starts from the fiber over $\delta_{K_{j}}(0)=-S$, but the Lagrangian is only over the part of the arc for $t>1$ because in $Y^{0}$ we puncture the base at 0 which is where the curve passes through at $t=1$. In other words, $W_{0}(K_{j})$ in the cylinder model is a cotangent fiber of $T^{*}S^{1}$ at height $\theta=\theta_{h}(K_{j})$.
@@
3.2. Morphism groups
3.2.1. Wrapping conventions in the cochain complex
For two transversely intersecting Lagrangians $L_{0},L_{1}$, the Floer cochain complex is

### v2

@@ 1.2. Statement of main theorem
There exist two full subcategories $\mathcal{A}_{L}$ (Lagrangians $L$ fibered over U-shapes) and $\mathcal{A}_{K}$ (Lagrangians $K$ fibered over cotangent fibers) of the Fukaya category $H^{*}FS(Y^{0},W_{0})$ which are equivalent to $D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$ and $D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$, respectively. Furthermore, for the full subcategory $\mathcal{A}_{L,K}:=\mathcal{A}_{L}\cup\mathcal{A}_{K}$, there is an equivalence
(1.11) $\mathcal{A}_{L,K}\xrightarrow[]{\cong}D^{b}_{\mathcal{L}}\mathrm{Coh}(X).$
@@
Note that $D^{b}\text{Coh}_{\mathcal{L}}(V)$ is referred to as $\mathcal{B}_{\mathcal{L}}$ in [ACLL_tori, §1.2.2].
@@
We can generalize the definition of $X$ and $(Y^{0},W_{0})$ by taking $H$ to be a theta divisor in a principally polarized abelian variety $V_{\tau}$ of any dimension, for a generic parameter $\tau$ in the Siegel upper half space. We expect the results of this paper can then be generalized to HMS between $X$ and $(Y^{0},W_{0})$, assuming HMS for theta divisors. This will involve generalizing the symplectic form in [Ca20], including a B-field, equipping Lagrangians with connection 1-forms with curvature determined by the B-field, replacing $\frac{i}{2\pi}\log t\left(\begin{matrix}2&1\\ 1&2\end{matrix}\right)$ with a general $\tau$, and generalizing calculations from $g=2$ here for an abelian surface to $g$ for an abelian variety of arbitrary dimension.
@@
In this paper, we assume trivial local systems on the Lagrangians as in [Ca20]. Global HMS for principally polarized abelian varieties of any dimension and with the most general connection 1-forms on Lagrangians is proved on expected generating sets in [ACLL_tori]. Here and in [Ca20], we consider fiber Lagrangians $\ell_{k,[0]}$ and $\omega_{B}=0$ in the notation of [ACLL_tori, §1.2.3 and Equation (2.36)].
Remark 1.11.
The Lagrangians in $\mathcal{A}_{L}$ and $\mathcal{A}_{K}$ do not bound discs. To define $FS(Y^{0},W_{0})$ one must account for the Lagrangians that do, with bounding cochains as in [fooo1, §3] for compact symplectic manifolds without superpotential. For a suitable definition of $FS(Y^{0},W_{0})$ with bounding cochains, upgrading to an $A_{\infty}$-result, and for a suitable adaptation of core HMS [PS23], we expect this theorem can be upgraded to the originally conjectured equivalence in Equation (1.1) which would be $D^{b}\text{Coh}(X)\cong D^{\pi}FS(Y^{0},W_{0})$.
Remark 1.12.
We directly construct Lagrangians and calculate their morphisms in $(Y^{0},W_{0})$. Another approach could be to combine the left diagram of Figure 1 for $H^{0}FS(Y,-v_{0})$ from [Ca20], prove a similar diagram for $H^{0}FS(V^{\vee}\times\mathbb{C}^{*}_{y^{\vee}},y^{\vee})$ as in the right diagram of Figure 1 where $V^{\vee}$ is the SYZ mirror of $V$, and then combine them as suggested in [Aur18, Remark 1.5] and [sylvan_orlov, §3.3: Gluing].
$D^{b}_{\mathcal{L}}\mathrm{Coh}(V)$$D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$$H^{0}\mathrm{Fuk}(V^{\vee})$$H^{0}{FS}(Y,-v_{0})$$\scriptstyle{\iota^{*}}$$\scriptstyle{\text{HMS on }V}$$\scriptstyle{\text{HMS on }H=\Sigma_{2}}\scriptstyle{\cup}$
$D^{b}_{\mathcal{L}}\mathrm{Coh}(V)$$D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$$H^{0}\mathrm{Fuk}(V^{\vee})$$H^{0}{FS}(V^{\vee}\times\mathbb{C}^{*}_{y^{\vee}},y^{\vee})$$\scriptstyle{\mathrm{pr}_{V}^{*}}$$\scriptstyle{\text{HMS on }V}\scriptstyle{\cup}$
@@
Outline of the paper. In Section 2 we describe the B-side and the semi-orthogonality of $D^{b}\text{Coh}(X)$. In Section 3 we describe the A-side. In Subsection 3.1 we define the mirror Lagrangian objects. In Subsection 3.2 we define the morphisms in the Fukaya subcategories of $(Y^{0},W_{0})$ by showing the continuation maps are well-defined. We define how Lagrangians wrap in §3.2.1 and outline the rest of Subsection 3.2 with an overview of how to pass from the cochain complex to the localized category in §3.2.2. Then in §3.2.3 we define the quasi-units we localize at, and in §3.2.4 we define their inverses, the continuation maps. Finally in §3.2.5 we define the morphisms by taking a homotopy colimit over multiplication by quasi-units, using a model from [AA24] that amounts to wrapping one way on the input Lagrangian and in the opposite direction on the output, in Equation (3.42). In Subsection 3.3 we define the Lagrangian gradings, which are part of the data of a D-brane in the Fukaya category. Then in Subsection 3.4 we calculate morphisms in Lemmas 3.25 (within $\mathcal{A}_{L}$), 3.27 ($\mathcal{A}_{L}$ to $\mathcal{A}_{K}$), 3.28 ($\mathcal{A}_{K}$ to $\mathcal{A}_{L}$), and 3.29 (within $\mathcal{A}_{K}$). In Subsection 3.5 we calculate the product structure on both sides. This requires an analysis of gradings to determine what intersection points appear in the product, which is Lemma 3.31. We then prove the main theorem, Theorem 1.7, in Section 4.
Acknowledgements. We thank Denis Auroux for helpful comments in formulating the direction of this project and for explaining his joint work [AA24]. We thank Hiro Lee Tanaka, Sheel Ganatra, and Zack Sylvan for helpful discussions on wrapping in Fukaya categories, Christian Schnell and Alexander Kuznetsov for helpful discussions on semi-orthogonal decompositions of $D^{b}\text{Coh}$, and Katrin Wehrheim for helpful discussions on moduli spaces. Thanks to Haniya Azam, Hiro Lee Tanaka, and Sheel Ganatra for helpful comments on an earlier draft.
@@ 2. B-side semi-orthogonal decomposition
2.1. Morphisms from $V\times\mathbb{C}$ to $H$
@@
(2.1) $\operatorname{Ext}(\pi^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}}),j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{p}(-1)))=0.$
2.2. Morphisms from $H$ to $V\times\mathbb{C}$
By the adjunction formula, $K_{X}=\pi^{*}K_{V\times\mathbb{C}}\otimes\mathcal{O}(E)$. As in [Huy05, Proposition 2.5.5], we define the relative canonical bundle as follows,
@@ 2.2. Morphisms from $H$ to $V\times\mathbb{C}$
where $\mathcal{O}_{E}(E)=j^{*}\mathcal{O}(E)=\mathcal{O}_{p}(-1)$. Furthermore, $j^{!}$ admits a left adjoint $j_{*}$ for $j:E\to X$ where
@@
Note that $\otimes\mathcal{O}_{E}(E)$ is an auto-equivalence of $D^{b}\text{Coh}(X)$, thus its contribution cancels. Also, the statement that Equation (1.9) commutes is that $\pi\circ j=i\circ p$. Thirdly, $p:E\to H\times\{0\}$ is surjective and $p^{*}$ is fully faithful, an isomorphism on hom sets. Hence we can simplify to obtain
@@
3. A-side semi-orthogonal decomposition
The Fukaya category roughly consists of Lagrangian submanifolds as objects and cochain complexes generated by their intersection points as morphisms, graded by a Maslov index. The differential counts bigons between input and output intersection points, weighted by area (and also holonomy around the boundary of the disc, which we do not consider here, see Remark 1.10). Hamiltonian-equivalent Lagrangians should be isomorphic in the derived Fukaya category. This has led to increasingly more general definitions of the Fukaya category, as more mirrors are constructed.
In the Fukaya category of a closed symplectic manifold (compact with no boundary), no wrapping is needed for closed Lagrangians. This is the original $A_{\infty}$-category [fooo1]. For a non-closed symplectic manifold, wrapping is needed for non-closed Lagrangians. So non-compact manifolds and those with boundary have wrapping. For example, the Fukaya category in [AS10, GPS20], which wraps with stops, applies to non-compact exact Liouville manifolds. [GPS24, §3.3] describe wrapping at a puncture as a correspondence between stop removal and localization at continuation maps.
Non-compact manifolds that appear as SYZ mirrors in many known examples are fibrations over $\mathbb{C}$ or over a subset with boundary; these are the LG models. Seidel [seidel] pioneered the first constructions of their Fukaya categories, specifically for non-compact exact Lefschetz fibrations. Therefore more generally for a symplectic fibration LG model, we refer to its Fukaya category as the Fukaya-Seidel category. To equip an LG model with a Fukaya-Seidel category, the base and the fiber determine the wrapping needed. For example, fibers given by closed manifolds have no wrapping, $(\mathbb{C}^{*})^{n}$-fibers have wrapping, the base $\mathbb{C}$ of an LG model wraps but not fully due to one stop, and punctures in the base or fiber introduce full wrapping around the puncture. Once the wrapping behavior is determined, the Fukaya-Seidel category of the symplectic fibration LG model can be constructed by defining the continuation maps and showing they are well-defined.
@@ 3. A-side semi-orthogonal decomposition
In [AA24] a fiberwise wrapped Fukaya-Seidel category is defined for non-compact non-exact LG models. In our LG model $(Y^{0},W_{0})$, there is one puncture in the base and generic fibers are closed $T^{4}$. Since the fibration is not Lefschetz, we avoid the singular fiber. So the projection under $W_{0}$ of Lagrangians in $\mathcal{A}_{L}$ are U-shapes around the critical value. And in $\mathcal{A}_{K}$ they project under $W_{0}$ to a curve that can fully wrap in the base around the puncture. All projected Lagrangians still wrap partially at infinity due to the one stop for LG models [ushape, Appendix A], [GPS20, §1.2, Example 1], [AA24]. Our continuation maps are defined in Definition 3.9.
Note that the complex dimension of the generalized SYZ mirror $(Y,-v_{0})$ to $\Sigma_{2}$ is two larger than that of $\Sigma_{2}$; if one takes $\text{Crit}(v_{0})$ as the mirror, which is complex dimension 1, then a Fukaya category on that has been constructed in [AEK24].
3.1. Objects in $\mathcal{A}_{L}$, $\mathcal{A}_{K}$
@@
The base of $W_{0}$ is $\mathbb{C}^{*}$ which is topologically a cylinder, and a generic fiber is $T^{4}$. So there are two pictorial representations for the base: $\mathbb{C}^{*}$ and the cylinder. Curves drawn in a plane represent projections of Lagrangians under $W_{0}$, while curves drawn on a cylinder represent projections of Lagrangians under $W_{0}$ composed with $\mathbb{C}^{*}\to\mathbb{R}\times S^{1}$ given by $e^{2\pi(\mu+i\theta)}\mapsto(\mu,\theta)$. Lagrangians are “thimble-like” in the sense of [GPS24, Definition 8.19].
@@
Definition 3.1 (Parallel transport, [symp_intro, Equation (6.2)]).
Let $\Phi_{\gamma(0)\to\gamma(t)}:W_{0}^{-1}(\gamma(0))\to W_{0}^{-1}(\gamma(t))$ denote parallel transport in $(Y^{0},W_{0})$. It is taken with respect to $\omega|_{Y^{0}}$, for $\omega$ on $Y$ defined in [Ca20, §3.5], over an embedded path $\gamma:[0,t]\to\mathbb{C}^{*}$ avoiding the critical value $\text{critv}(W_{0})=-T^{\epsilon}$.
We define two subcategories $\mathcal{A}_{L},\mathcal{A}_{K}$ of the Fukaya-Seidel category of $(Y^{0},W_{0})$. Define a point
@@
where $(\theta_{1},\theta_{2})=\frac{1}{2\pi}(\arg(t_{1}),\arg(t_{2}))$ for $(t_{1},t_{2})\in(\mathbb{C}^{*})^{2}$ the affine coordinates corresponding to monomials $\chi^{(1,0,0)}$ and $\chi^{(0,1,0)}$ on the polytope $\Delta_{Y}$ of Equation (1.6), and $(\xi_{1},\xi_{2})$ are defined in Equation (1.4).
Definition 3.3 (Objects of $\mathcal{A}_{L}$).
Objects of $\mathcal{A}_{L}$ are $\{L_{j}[k]\}_{j,k\in\mathbb{Z}}$ defined as
@@
such that $r(t),\theta(t):\mathbb{R}\to\mathbb{R}$ are smooth with
(3.2) $(r(0),\theta(0))=(S,-\frac{1}{2}),\qquad r(t)\propto|t|,\dot{\theta}(t)=0\;\;\forall t\notin r^{-1}([R_{2},R_{4}]).$
Let $\gamma_{L_{j}}(t)$ initially near $t=0$ stay within rings $R_{2}$ and $R_{4}$ in Figure 3, and take a U-shape around the critical value at $\times=-T^{\epsilon}$ and below the puncture at 0. In particular, $\gamma_{L_{j}}(0)=-S$. When $r(t)$ reaches $R_{4}$, the ends are at constant angles away from $\pi$
(3.3) $\theta(t)=\begin{cases}\theta_{h_{1}}(L_{j})&\forall t\gg 0\\ \theta_{h_{2}}(L_{j})&\forall t\ll 0\end{cases}$
and strictly increasing radial direction. See $W_{0}(L_{j_{1}})$ in Figure 8, where the $\theta_{h_{1}}(L_{j})$ end corresponds to the lower end of the U-shape.
@@
U-shaped Lagrangians have appeared in the literature such as [ushape, §A.2], [ACLL_LG], and [AA24]. Here we use the model that ends of the U-shape project to angular rays away from 0 in $\mathbb{C}^{*}$.
Definition 3.5 (Objects of $\mathcal{A}_{K}$).
Objects of $\mathcal{A}_{K}$ are $\{K_{j}[k]\}_{j,k\in\mathbb{Z}}$ defined as
@@
such that $r(t),\theta(t):[0,\infty)\to\mathbb{R}$ are smooth with
(3.4) $(r(0),\theta(0))=(S,-\frac{1}{2}),\quad r(1)=0,\quad(r(t),\theta(t))=(t-1,\theta_{h}(K_{j}))\;\;\forall t>1$
so that $\delta_{K_{j}}$ is an arc from S-S to 0 avoiding the critical value at $T^{\epsilon}$ by going above it, and is then a single arc at fixed angle $\theta_{h}(K_{j})$. To be the correct mirror, we take the arc to lie above any U-shape ends in the $\mathbb{C}^{*}$ model, so it lies counterclockwise from the U-shape ends at angles $\theta_{h_{1}}(L_{j^{\prime}}),\theta_{h_{2}}(L_{j^{\prime}})$ for all $j^{\prime}\in\mathbb{Z}$, and clockwise from the ray at $\pi$. Parallel transport starts from the fiber over $\delta_{K_{j}}(0)=-S$, but the Lagrangian is defined only over the part of the arc for $t>1$ because in $Y^{0}$ we puncture the base at 0, which is where the curve passes through at $t=1$. See $W_{0}(K_{j})$ in Figure 6(b). In the cylinder model, $W_{0}(K_{j})$ is a cotangent fiber of $T^{*}S^{1}$ at height $\theta=\theta_{h}(K_{j})$, see $W_{0}(K_{j})$ in Figure 6(a).
@@
Remark 3.6.
We fix the arc to go above the U-shaped ends when viewed in the base of $W_{0}$. Had we taken it to go below, it would swap the order of the semi-orthogonal decomposition and be mirror to a different embedding of the components on the B-side too.
3.2. Definition of morphisms
In this section we define the morphisms. We give some motivation for the process. The morphisms are given by the wrapped Lagrangian intersection Floer cohomology theory, described on the chain-level in §3.2.1 as intersections of Lagrangians and denoted by $CF(L_{0},L_{1})$. Furthermore, Hamiltonian-isotopic Lagrangians should be quasi-isomorphic, meaning isomorphic when we derive the category. There is a convenient tool from homotopy theory that does the job for non-compact Lagrangians, called a homotopy colimit [AS10, AA24]. As explained in the introduction to [Hi03], a goal of homotopy theory is to identify homotopic maps, but a homotopy on topological spaces may not make sense in other categories. To identify homotopic morphisms $f$ and $g$, we define a set called the weak equivalences to be where fgf-g lands in, and then localize the category at weak equivalences. To take a homotopy colimit of a diagram one needs a model category: a category with three sets of morphisms called weak equivalences, fibrations, and cofibrations. Quillen axiomatized such a category as a place where one can do homotopy theory, calling them “a category of models for a homotopy theory” [Qu67]. Note that more modern approaches in homotopy theory use $\infty$-categories instead of model categories. In the wrapped Fukaya category, weak equivalences will be multiplication by quasi-units. Furthermore, we will take the images of objects in the Fukaya category under the Yoneda embedding, and work inside the resulting triangulated category.
We motivate how to glue together equivalent objects, that is, how to identify homotopic morphisms. An example [We94, §1.5] is that of making a simplicial complex $X$ contractible by gluing it to a point. Taking its cone $CX$ then $C_{j}(CX)=C_{j}(X)\oplus C_{j-1}(X),\;j\geq 1$ from simplices fully in the base and those through the cone vertex $s$. Checking the differential and quotienting by $C_{0}(s)$ generated by the cone vertex, this is just $C_{j}(CX)/C_{j}(s)=\text{Cone}_{j}(\mathbf{1}_{\bullet}:C_{\bullet}(X)\to C_{\bullet}(X))$. So the notion of topological and algebraic cone coincide.
In practice when computing morphisms of Lagrangians in a Landau-Ginzburg model, the informal way of doing so is to “push the ends of the input Lagrangian, when projected to the base, to lie above the ends of the output Lagrangian.” In other words, if $L_{0}^{+}$ is $L_{0}$ with the ends pushed up, we would like a morphism $L_{0}\to L_{1}$ to be a roof $L_{0}^{+}$$L_{0}$$L_{1}$$\scriptstyle{e}$$\scriptstyle{p}$ where $e\in CF(L_{0}^{+},L_{0})$ is a quasi-unit (defined in Equation (3.22)) and $p\in L_{0}^{+}\cap L_{1}$ is an intersection point, a morphism of $CF(L_{0}^{+},L_{1})$. This roof is the standard construction of a left fraction [We94, Prototype 10.3.5], a morphism in a category localized at $e$. To compose roofs would involve a pullback along $L_{0}^{+}\rightarrow L_{1}\leftarrow L_{1}^{+}$ [We94, Definition 10.3.4, Ore condition]. However, it is not evident what this pullback would be in practice.
We instead take a homotopy colimit over Yoneda modules $\mathcal{Y}_{L^{k}}=\mathcal{O}(-,L^{k})$ in the abelian category of $\mathcal{O}$-modules for a category $\mathcal{O}$ directed by amount wrapped, defined in Equation (3.20). The objects $L^{k}$ are indexed by wrapping an amount $k$ defined in Equation (3.17). The homotopy colimit identifies objects that have a homotopy between them, preserving weak equivalences which are when a map of spaces induces an isomorphism on homotopy groups (so a particular inverse isn’t needed, as for a strong equivalence). This is similar to how a colimit glues along connecting maps. However we need the homotopy version, for example a limit such as the pullback of $0\hookrightarrow[0,1]\hookleftarrow 1$ is empty even though $0,1,[0,1]$ are all homotopy equivalent, while the homotopy limit would be $\{\gamma:[0,1]\to[0,1]\mid\gamma(0)=0,\gamma(1)=1\}$, [Ri14, Exercise 6.5.4]. Weak equivalences here are multiplcation by quasi-units, and we localize at quasi-units in order to identify non-compact Lagrangians equivalent under wrapping in the derived Fukaya-Seidel category.
We follow [AA24, §3.5], where they use standard constructions similar to [AS10, §(3g)], [GPS20, §3.1.3]. These refer to [LO06] for localization of $A_{\infty}$-categories, a process that generalizes localization of categories [We94, §10.3], [Ne01, §1], [Ri14, Example 11.5.11].
3.2.1. Wrapping conventions and the cochain complex
For two transversely intersecting Lagrangians $L_{0},L_{1}$, the Floer cochain complex without local systems is